# ReviewFlow: Intelligent Scaffolding to Support Academic Peer Reviewing

### 05 Feb 2024 v1

@@ 7.1. What role should machine intelligence play in this review process?
Prior work indicates that well-design scaffolding can help novices operate more like experts (Yuan et al., 2016). The concept of Zone of Proximal Development (ZPD) represents the space between what a learner is capable of doing unsupported and what the learner cannot do even with support (Harland, 2003; Wass et al., 2011). Scaffolding works most effectively when it meets novices in their ZPD, meeting people where they need support. and tapering off when they have internalized the best practices. The scaffolding in ReviewFlow is intelligent and contextual. For instance, users get reflection cues that adapt specifically to the current section, however the scaffolding does not adapt to the user. Future work should focus on making scaffolding more interactive and adaptive to a learner’s current level of knowledge.
@@ 7.2. How can intelligent scaffolding support novices across the science ecosystem?
Our study of ReviewFlow indicates novice reviewers can benefit from intelligent scaffolding: it helps them evaluate the submission and write more comprehensive reviews. Participants mentioned how they gradually learned about the research community’s practices and became more confident in their review writing. Beyond just writing reviews, the strategies for intelligent scaffolding built into ReviewFlow have the potential to provide value to other aspects of the peer review ecosystem (Shah et al., 2018). For example, similar intelligent scaffolding can be used to support novice reviewers to revise their papers and write a persuasive rebuttal. The process of meta-reviewing, or summarizing independent reviews, could also benefit from scaffolding since there are new people jumping into that role each cycle.
Future deployments of this type of intelligent scaffolding would require careful consideration and round-table discussions within the research community. Previous research revealed that writing with an opinionated language model can affect participants’ opinions in writing and attitude towards social topics (Jakesch et al., 2023). In our study, even though our scaffolding was carefully designed to allow the user to drive the process and to avoid biasing the decision, we still heard participants who were concerned about other people using such a system. Interestingly, very few participants were concerned that their own decision or writing was being biased by the AI, instead, they worried about how others would appropriate the intelligent support and how it might erode a community’s trust in the peer review system. To deploy the system in the future and build trust with people in the community, we need to make sure that the system is robust and trustworthy (Wang et al., 2023). A real-world deployment would need to have a large-scale consent process and commitment from key stakeholders.
Beyond its application in peer reviewing, intelligent scaffolding could support a range of of complex knowledge work in science, such as paper reading, literature reviewing, sensemaking, and paper writing (Palani et al., 2023; Liu et al., 2023a; Zhang et al., 2023; Lin et al., 2024). Notable projects like the semantic reader project have developed interactive and dynamic reading interfaces to aid paper reading and citation discovery (Lo et al., 2023; Kang et al., 2022; August et al., 2022; Fok et al., 2023). PaperPlain introduced key questions for readers to consider, highlighting corresponding answers within the paper content (August et al., 2022). These studies suggest broader possibilities for incorporating intelligent, process-driven scaffolding into science work.
@@ 7.3. Ethical considerations of AI scaffolding for academic review
Generative AI and LLMs introduce numerous ethical considerations in the design of human-AI collaboration systems. In the context of academic publishing and peer review, a primary concern involves the potential violation of academic integrity and harm to authorship when directly incorporating automatically generated content into writing artifacts. We try to mitigate this in ReviewFlow by constraining the use of LLMs to primarily sensemaking activities, like guided note taking. In the final step, ReviewFlow synthesizes an outline, not paragraphs, even though LLMs are quite capable of doing so. To limit data sharing, ReviewFlow only sends the LLM the reviewer’s notes and highlighted segments, not the full paper. While individual reviewers may still choose to (mis)use LLMs in these ways, ReviewFlow subtly prioritizes learning and thinking over getting the job done fast.
Another concern revolves around the tendency of LLMs to create inaccurate information or to misled people(Huang et al., 2019; Jakesch et al., 2023). Inline with our scaffolding strategy, ReviewFlow includes a checklist popup that allows users to engage in self-reflection, with reminders to proofread and fact-check their review. To enhance transparency in the writing process, future systems might explore the development of intelligent highlighting. For example, highlights could directly color-code the portions that were user-generated and those that were automatically generated, providing a clear visual indication and promoting a more transparent and accountable writing process.
@@ 8. Limitations
Our study has several limitations. First, ReviewFlow combined multiple scaffolding strategies into one tool, leaving as future work to understand to what extent each strategy impacted the outcomes. Ablation studies could provide more insight on the effectiveness of each scaffold. Second, we simulated a mock peer review scenario where users had about one hour to read and write a review for a short paper. In practice, as we learned in our preliminary interviews, writing peer reviews can hours. A longer study could give insights on the enduring value of scaffolding. Third, we selected two HCI papers to use in our experiment, but different research domains and communities have different guidelines and standards for reviewing. Deploying the system across research communities may bring new insights into the generalization of scaffolding strategies. Lastly, the underlying machine models will evolve over time, which can impact the performance of ReviewFlow, for better or worse. Continuous updates and adaptations to the latest AI models would be required to maintain the tool’s effectiveness and relevance.
@@ 9. Conclusion
This research explore techniques for integrating LLMs into intelligent scaffolding for academic peer reviews. Our formative studies found that expert reviewers adopted a workflow of annotating, note-taking, and synthesizing notes before writing, while novices lacked perspective on the prior work in the domain and reviewing standards. Modeling the expert workflow, we developed ReviewFlow –LLM-supported workflow that scaffolds novices using contextual reflection cues, in-situ knowledge support, and notes-to-outline synthesis. In a within-subject experiment with 16 novice reviewers, we found that ReviewFlow led to more comprehensive peer review and higher self-efficacy on the task. We further discuss the implication of intelligent scaffolding in knowledge work.

### 26 Feb 2024 v2

@@ 7.1. What role should machine intelligence play in this review process?
Prior work indicates that well-design scaffolding can help novices operate more like experts (Yuan et al., 2016). The concept of Zone of Proximal Development (ZPD) represents the space between what a learner is capable of doing unsupported and what the learner cannot do even with support (Harland, 2003; Wass et al., 2011). Scaffolding works most effectively when it meets novices when they need support and tapering off when they have internalized the best practices. The scaffolding in ReviewFlow is intelligent and contextual. For instance, users get reflection cues that adapt specifically to the current section.
@@ 7.2. How can intelligent scaffolding support novices across the science ecosystem?
Our study of ReviewFlow indicates novice reviewers can benefit from intelligent scaffolding: it helps them evaluate the submission and write more comprehensive reviews. Participants mentioned how they gradually learned about the research community’s practices and became more confident in their review writing. Beyond just writing reviews, the strategies for intelligent scaffolding built into ReviewFlow have the potential to provide value to other aspects of the peer review ecosystem (Shah et al., 2018). For example, similar intelligent scaffolding can be used to support novice reviewers to revise their papers and write a persuasive rebuttal. The process of meta-reviewing, or summarizing independent reviews, could also benefit from scaffolding since there are new people jumping into that role each cycle (Sun et al., 2024).
Future deployments of this type of intelligent scaffolding would require careful consideration and round-table discussions within the research community. Previous research revealed that writing with an opinionated language model can affect participants’ attitudes towards social topics in writing (Jakesch et al., 2023). In our study, even though our scaffolding was carefully designed to allow the user to drive the process and to avoid biasing the decision, we still heard participants who were concerned about other people using such a system. Interestingly, very few participants were concerned that their own decision or writing was being biased by the AI, instead, they worried about how others would appropriate the intelligent support and how it might erode a community’s trust in the peer review system. To deploy the system in the future and build trust with people in the community, we need to make sure that the system is robust and trustworthy (Wang et al., 2023). A real-world deployment would need to have a large-scale consent process and commitment from key stakeholders.
Beyond its application in peer reviewing, intelligent scaffolding could support a range of of complex knowledge work in science, such as paper reading, literature reviewing, sense-making, and paper writing (Palani et al., 2023; Liu et al., 2023a; Zhang et al., 2023; Lin et al., 2024). Notable projects like the semantic reader project have developed interactive and dynamic reading interfaces to aid paper reading and citation discovery (Lo et al., 2023; Kang et al., 2022; August et al., 2022; Fok et al., 2023). These studies suggest broader possibilities for incorporating intelligent, process-driven scaffolding into science work.
@@ 7.3. Ethical considerations of AI scaffolding for academic review
Generative AI and LLMs introduce numerous ethical considerations in the design of human-AI collaboration systems. In the context of academic publishing and peer review, a primary concern involves the potential violation of academic integrity and harm to authorship when directly incorporating automatically generated content into writing artifacts. We try to mitigate this in ReviewFlow by constraining the use of LLMs to primarily sensemaking activities, like guided note-taking. In the final step, ReviewFlow synthesizes an outline, not paragraphs, even though LLMs are quite capable of doing so. To limit data sharing, ReviewFlow only sends the LLM the reviewer’s notes and highlighted segments, not the full paper. While individual reviewers may still choose to (mis)use LLMs in these ways, ReviewFlow subtly prioritizes learning and thinking over getting the job done fast.
Another concern revolves around the tendency of LLMs to create inaccurate information or to mislead people(Huang et al., 2019; Jakesch et al., 2023). In line with our scaffolding strategy, ReviewFlow includes a checklist pop-up that allows users to engage in self-reflection, with reminders to proofread and fact-check their review. To enhance transparency in the writing process, future systems might explore the development of intelligent highlighting. For example, highlights could directly color-code the portions that were user-generated and those that were automatically generated, providing a clear visual indication and promoting a more transparent and accountable writing process.
@@ 8. Limitations
Our study has several limitations. First, ReviewFlow combined multiple scaffolding strategies into one tool, leaving future work to understand to what extent each strategy impacted the outcomes. Ablation studies could provide more insight into the effectiveness of each scaffold. For instance, a study with three experimental conditions, each providing only one type of scaffolding would allow for a more comprehensive evaluation. Second, we simulated a mock peer review scenario where users had about one hour to read and write a review for a short paper. In practice, as we learned in our preliminary interviews, writing peer reviews takes hours. A longer study could give insights into the enduring value of scaffolding on a longer paper. Third, we selected two papers in HCI to use in our experiment, but different research domains and communities have different guidelines and standards for reviewing. Deploying the system across research communities may bring new insights into the generalization of scaffolding strategies. Lastly, the underlying machine models and LLMs will keep improving, which can impact the performance of ReviewFlow, for better or worse. Continuous updates and adaptations to the latest AI models would be required to maintain the tool’s effectiveness and relevance.
@@ 9. Conclusion
This research explores techniques for integrating LLMs into intelligent scaffolding for academic peer reviews. Our formative studies found that expert reviewers adopted a workflow of annotating, note-taking, and synthesizing notes before writing, while novices lacked perspective on the prior work in the domain and reviewing standards. Modeling the expert workflow, we developed ReviewFlow –LLM-supported workflow that scaffolds novices using contextual reflection cues, in-situ knowledge support, and notes-to-outline synthesis. In a within-subject experiment with 16 novice reviewers, we found that ReviewFlow led to more comprehensive peer review and higher self-efficacy on the task. We further discuss the implication of intelligent scaffolding in knowledge work.
Appendix
Features Prompt Examples
Section-level contextual cues
System role: You are an expert peer reviewer providing guided questions for novice peer reviewers to help them write better reviews.
Importance: Is the problem addressed in the submission, concerning the video explicitness of a video conferencing interface for virtual
Context: Given this paper’s abstract: { abstract }
studying, important to the CSCW community?
Template: What could be a helpful guided question for novice reviewers to evaluate the { aspect } of this paper for this section of the paper? Please output 1 concise question with a maximum of 25 words
Novelty: How does this study relate to existing work in the field of virtual social activities and video streaming? Does it introduce a novel approach?
phrase-level contextual cues
System role: You are an expert peer reviewer providing guided questions for novice peer reviewers to help them write better reviews.
Validity: Does the Method section provide clear and well-justified explanations of the research prototype’s design, including the choice of video
Context: Given this paper’s abstract: { abstract }
versions and activity recognition method?
Template: What could be a helpful guided question for novice reviewers to evaluate the { aspect } of this paper for this specific paragraph:{ paragraph }. Please output 1 concise question with a maximum of 25 words for this specific paragraph.
Clarity: How comprehensive does the method explain the recruitment process, demographics, and procedures of the user study, ensuring clarity for a CSCW audience?
Notes-to-outline synthesis
System role: I am building a web application that helps novice peer reviewers write better peer reviewers. Your task is to synthesize the reviewer’s notes into an outline. Please provide an outline with sections of the summary of the paper, strengths, and weaknesses. To accomplish this task, I am providing an abstract of the reviewed paper, each of reviewer’s notes and corresponding contents of the reviewed paper for context, and the topics that should be used to generate the weakness part of the outline.
Summary:     Explored new design strategies for virtual studying focusing on sharing presence.     Context of the study is the advent of COVID-19 and arising of new virtual social activities.     Investigated minimizing video explicitness in virtual studying using various strategies.     The study involved four virtual studying teams that used prototypes and participated in interviews.
Template: Here is the paper abstract:{abstract}; Here are the user annotations that contain Note and paper content only as context:{notes}. Please create three to five concise bullet points for each section (strengths, weaknesses) out of the reviewer’s notes for the templated outline. The weakness part of the outline should be based on the topics for weakness but with detailed descriptions using only the notes under each topic. The topics should have a length of at most 10 words. The details should have a least 10 words and be a complete sentence.
Strength:     Relevance to current societal changes due to COVID-19. – Paper timely addresses the shift to virtual activities sparked by COVID-19.     Addresses an emerging need in virtual collaborative environments. – Investigates novel video conferencing interfaces for enhanced privacy in virtual studying. Weakness:     Limited scope of the study sample. – Sample size is limited to four virtual studying teams, restricting broader applicability of findings.     Lack of precision in relating strategies to prior research. – The paper lacks precise connections between its strategies and the existing research, undermining its innovative claim.     Insufficient citations on related work, specifically Barbara’s contributions – Needs more references to Barbara’s work to establish the paper’s context.
Table 3. Prompts and examples used in the ReviewFlow system.